# The WMDP Benchmark: Measuring and Reducing Malicious Use With Unlearning

### v2

@@ 5.2 Quantitative Evaluation
To assess the efficacy of the methods, we examine the forget performance and retain performance of the unlearned models. We see that Cut is able to unlearn WMDP-Bio and WMDP-Cyber while maintaining performance on MMLU (Section 5).Forget performance.
We measure forget performance by evaluating the knowledge of models on WMDP with both question-answering (QA) and probing. QA evaluation. In the future, LLMs may be used by adversaries as knowledge engines for developing weapons. Under an API-access threat model, adversaries only receive output tokens and logits, without access to internal activations. Hence, we evaluate the QA accuracy of models on WMDP. We use a zero-shot question-answer format (Section B.2), taking the top logit between A, B, C, and D as the answer choice. For comparison, we also benchmark GPT-4 zero-shot on each of these tasks. As language models are sensitive to the prompting scheme (Sclar et al., 2023), we use lm-evaluation-harness (Gao et al., 2021) to standardize prompts. QA results. We assess whether Cut is able to reduce QA accuracy on WMDP in Table 1. For both zephyr-7b and Yi-34b, Cut is able to drop performance to near random accuracy on WMDP-Bio and WMDP-Cyber, while other baselines struggle to drop accuracy on WMDP-Bio and WMDP-Cyber without crippling model performance on MMLU. We provide a more comprehensive table of results in 1. Probing evaluation. While evaluating QA accuracy measures the primary risk of the API-access threat model, it fails to assess whether knowledge has been fully removed from the models. Models may possess more knowledge than is revealed in their output logits (Burns et al., 2022); for instance, the unlearned model may still retain hazardous knowledge, but refuse to answer. Thus, we test whether unlearned models can be probed to recall unlearned information. We train a 4-way linear probe on the unlearned Cut models. We use half of WMDP-Bio and WMDP-Cyber for training and hold out the other half for evaluation. We apply probing and report results for all layers of the model. Probing results. We assess whether probes are able to recover knowledge from a model unlearned with Cut in Figure 9. Across both categories and model sizes, linear probing only achieves slightly better than random accuracy. Linear probes are unable to extract unlearned information from the model, suggesting that Cut does not merely mask or hide the information superficially, but rather causes a substantial alteration that prevents the recall of the unlearned information.Retain performance.
@@
We measure the retain performance by evaluating models’ knowledge on MMLU and their fluency on MT-Bench. MMLU evaluation. To be practical, unlearning methods must maintain general knowledge while removing hazardous knowledge. To evaluate whether models retain general knowledge after unlearning, we reuse the earlier QA evaluation setup for MMLU. MMLU results. We report accuracy on subject-specific areas in MMLU (Figure 11). In contrast to other baselines which either fail to reduce performance on WMDP or greatly reduce performance on MMLU (Figure 11), Cut reduces performance on WMDP while maintaining overall MMLU accuracy. Moreover, Figure 11 shows that Cut retains performance on MMLU topics related to biology (college biology) and computer science (college CS), suggesting greater unlearning precision than the baselines. However, Cut greatly drops performance on the most similar topics to biosecurity (virology) and cybersecurity (computer security), suggesting the possibility for future work to improve retention of general capabilities during unlearning. As we use Wikitext as the retain set, Cut cannot determine exactly what knowledge to unlearn and retain. Thus, we encourage future work to employ our subject-specific biology and cyber retain sets (Section 4.2) to improve unlearning precision. Figure 10: Results across a hyperparameter search. Compared to the other baselines, Cut is most capable of reducing WMDP performance while maintaining accuracy on MMLU.
Figure 12: After applying Cut on Yi-34b, the GCG adversarial attack (Zou et al., 2023b) cannot extract hazardous knowledge within optimization steps, despite eliciting the same knowledge from base models in less than 50 steps.
@@ 5.3 Robustness Evaluation
A primary motivation for unlearning is ensuring that knowledge is irrecoverable, even when subject to optimization pressure. If unlearning is not resilient, the adversary can still jailbreak the model to access hazardous information after unlearning. We conduct a qualitative experiment using the GCG adversarial attack (Zou et al., 2023b) to measure whether dangerous knowledge is recoverable after performing Cut. We sample a single prompt from each of the WMDP-Bio and WMDP-Cyber datasets, slightly modify it such that the base Yi-34b models refuse to answer, and identify whether GCG can jailbreak the base and unlearned Yi-34b models to extract the correct answer (Section B.4). GCG can jailbreak the base Yi-34b models to answer these prompts in less than 50 gradient steps, while the unlearned models output gibberish even after steps, or over 7 hours of optimization on an NVIDIA A100 GPU (Figure 12). This is a signal towards the resilience of Cut, suggesting that unlearning persists even under optimization pressure.5.4 Generalization of WMDP to Hazardous Knowledge
@@
Figure 13: Unlearning on WMDP-Bio correlates with unlearning especially hazardous biology knowledge. This suggests that WMDP is a reasonable proxy measurement for hazardous knowledge.
@@
We follow the QA evaluation described in Section 5.2 and report the performance of zephyr-7b and Yi-34b before and after unlearning with Cut on this private set in Figure 13. Before and after unlearning, both models achieve similar accuracy on both the private set and WMDP. This result suggests WMDP is a reasonable proxy for especially hazardous knowledge.
@@ 6 Discussion
We discuss how unlearning on WMDP can tie in with other strategies to mitigate malicious use, such as structured API access. See Appendix D for a fuller discussion of the broader impacts of WMDP.6.1 How WMDP Mitigates Risk
@@
Unlearning on WMDP mitigates risk for both closed-source and open-source models. For closed-source models, unlearning reduces risk from malicious API finetuning (Zhan et al., 2023; Qi et al., 2023; Pelrine et al., 2023), as hazardous knowledge can be removed prior to serving the model. Furthermore, unlearning is a countermeasure against jailbreaks—even if they are jailbroken, unlearned models lack the knowledge necessary to empower malicious users (Figure 2). For open-source models, unlearning can expunge hazardous knowledge before such models are publicly released, limiting adversaries from repurposing open-source models out of the box. However, unlearning with on WMDP does not prevent relearning in open source models (Section B.5). We encourage future work towards mitigating risk in this pathway.6.2 Structured API Access
@@
WMDP complements the safety benefits of structured API access (Shevlane, 2022), where model developers provide an API for users to query and finetune models without full weight access. In this framework, ordinary users may query and finetune models with an API, but the model provider applies safety mechanisms, such as unlearning, prior to serving the model. However, approved users could obtain API access to the base model with full capabilities under strict guidelines, empowering the use of LLMs for benign or defensive applications while mitigating potential vectors of malicious use. For instance, OpenAI allows access of GPT-4 variants with fewer guardrails for red-teaming and biological malicious use experiments (OpenAI, 2023a, 2024). Structured access mitigates the concern that unlearning dual-use information will harm defenders. Structured access requires model developers to solve the “Know Your Customer” (KYC) challenge, which involves verifying the identity and intentions of customers before allowing them privileged interactions. For structured access, implementing KYC-like procedures can help mitigate the risks associated with malicious use by ensuring that only verified and trustworthy individuals or organizations are given the full capabilities of the model.7 Conclusion
@@
We propose a dataset, WMDP, to evaluate the potential of malicious use in LLMs. WMDP was developed by subject matter experts in biology, cybersecurity, and chemistry, and was filtered to remove sensitive or export-controlled information. Modern LLMs score highly on some aspects of WMDP, suggesting presence of hazardous knowledge. We propose machine unlearning as a safety intervention to reduce hazardous knowledge. Towards making progress on unlearning, we introduce Cut, an unlearning method that removes hazardous knowledge without significantly compromising general model performance. Cut also generalizes and successfully removes information from a private sensitive dataset. However, Cut reduces accuracy on closely related fields, such as introductory virology and computer security, demonstrating the need for continued research towards improved unlearning precision.Acknowledgments

### v3

@@ 5.2 Quantitative Evaluation
To assess the efficacy of the methods, we examine the forget performance and retain performance of the unlearned models. We see that RMU is able to unlearn WMDP-Bio and WMDP-Cyber while maintaining performance on MMLU (Section 5).
Forget performance.
We measure forget performance by evaluating the knowledge of models on WMDP with both question-answering (QA) and probing.
QA evaluation. In the future, LLMs may be used by adversaries as knowledge engines for developing weapons. Under an API-access threat model, adversaries only receive output tokens and logits, without access to internal activations. Hence, we evaluate the QA accuracy of models on WMDP. We use a zero-shot question-answer format (Section B.1), taking the top logit between A, B, C, and D as the answer choice. For comparison, we also benchmark GPT-4 zero-shot on each of these tasks. As language models are sensitive to the prompting scheme (Sclar et al., 2023), we use lm-evaluation-harness v0.4.2 (Gao et al., 2021) to standardize prompts.
QA results. We assess whether RMU is able to reduce QA accuracy on WMDP in Table 1. For both zephyr-7b and Yi-34b, RMU is able to drop performance to near random accuracy on WMDP-Bio and WMDP-Cyber, while other baselines struggle to drop accuracy on WMDP-Bio and WMDP-Cyber without crippling model performance on MMLU. We provide a more comprehensive table of results in 1.
Probing evaluation. While evaluating QA accuracy measures the primary risk of the API-access threat model, it fails to assess whether knowledge has been fully removed from the models. Models may possess more knowledge than is revealed in their output logits (Burns et al., 2022); for instance, the unlearned model may still retain hazardous knowledge, but refuse to answer. Thus, we test whether unlearned models can be probed to recall unlearned information. We train a 4-way linear probe on the unlearned RMU models. We use half of WMDP-Bio and WMDP-Cyber for training and hold out the other half for evaluation. We apply probing and report results for all layers of the model.
Probing results. We assess whether probes are able to recover knowledge from a model unlearned with RMU in Figure 9. Across both categories and model sizes, linear probing only achieves slightly better than random accuracy. Linear probes are unable to extract unlearned information from the model, suggesting that RMU does not merely mask or hide the information superficially, but rather causes a substantial alteration that prevents the recall of the unlearned information.
Retain performance.
We measure the retain performance by evaluating models’ knowledge on MMLU and their fluency on MT-Bench.
MMLU evaluation. To be practical, unlearning methods must maintain general knowledge while removing hazardous knowledge. To evaluate whether models retain general knowledge after unlearning, we reuse the earlier QA evaluation setup for MMLU.
@@
MMLU results. We report accuracy on subject-specific areas in MMLU (Figure 11). In contrast to other baselines which either fail to reduce performance on WMDP or greatly reduce performance on MMLU (Figure 11), RMU reduces performance on WMDP while maintaining overall MMLU accuracy. Moreover, Figure 11 shows that RMU retains performance on MMLU topics related to biology (college biology) and computer science (college CS), suggesting greater unlearning precision than the baselines. However, RMU greatly drops performance on the most similar topics to biosecurity (virology) and cybersecurity (computer security), suggesting the possibility for future work to improve retention of general capabilities during unlearning. As we use Wikitext as the retain set, RMU cannot determine exactly what knowledge to unlearn and retain. Thus, we encourage future work to employ our subject-specific biology and cyber retain sets (Section 4.2) to improve unlearning precision.
MT-Bench evaluation. Beyond retaining performance on academic multiple-choice questions, unlearned models should still maintain general conversational and assistant abilities. We evaluate RMU and all baselines on MT-Bench, a widely used metric for language model conversational fluency and helpfulness. We again evaluate GPT-4 as an upper bound for benchmark performance.
MT-Bench results. We report the MT-Bench performance of all models in Table 1. RMU roughly maintains performance on MT-Bench, with the score only decreasing on zephyr-7b and points on Yi-34b (out of a total possible of ). Because RMU still exhibits some degradation on MT-bench, particularly with Yi-34b, there is a need for further development of unlearning methods that can retain general assistant capabilities.
@@ 5.3 Robustness Evaluation
A primary motivation for unlearning is ensuring that knowledge is irrecoverable, even when subject to optimization pressure (Schwinn et al., 2024; Lynch et al., 2024). If unlearning is not resilient, the adversary can still jailbreak the model to access hazardous information after unlearning.
We conduct a qualitative experiment using the GCG adversarial attack (Zou et al., 2023b) to measure whether dangerous knowledge is recoverable after performing RMU. We sample a single prompt from each of the WMDP-Bio and WMDP-Cyber datasets, slightly modify it such that the base Yi-34b models refuse to answer, and identify whether GCG can jailbreak the base and unlearned Yi-34b models to extract the correct answer (Section B.3).
GCG can jailbreak the base Yi-34b models to answer these prompts in less than 50 gradient steps, while the unlearned models output gibberish even after steps, or over 7 hours of optimization on an NVIDIA A100 GPU (Figure 12). This is a signal towards the resilience of RMU, suggesting that unlearning persists even under optimization pressure.
Because we empirically only unlearn knowledge from three layers, RMU perhaps obfuscates knowledge more than unlearns it from the model. Thus, we investigate whether RMU is robust to finetuning in Appendix B.6. We emphasize, however, that finetuning after unlearning is not covered by our threat model, as closed-source LLM providers can always choose to apply unlearning immediately before model serving (Figure 2).
5.4 Generalization of WMDP to Hazardous Knowledge
@@
We follow the QA evaluation described in Section 5.2 and report the performance of zephyr-7b before and after unlearning with RMU on this private set in Figure 13. Before and after unlearning, both models achieve similar accuracy on both the private set and WMDP. This result suggests WMDP is a reasonable proxy for especially hazardous knowledge.
@@ 6 Discussion
We discuss how unlearning on WMDP can tie in with other strategies to mitigate malicious use, such as structured API access. See Appendix D for a fuller discussion of the broader impacts of WMDP.
6.1 How WMDP Mitigates Risk
Unlearning on WMDP mitigates risk for both closed-source and open-source models.
@@ 6.1 How WMDP Mitigates Risk
For closed-source models, unlearning reduces risk from malicious API finetuning (Zhan et al., 2023; Qi et al., 2023; Pelrine et al., 2023), as hazardous knowledge can be removed prior to serving the model. Furthermore, unlearning is a countermeasure against jailbreaks—even if they are jailbroken, unlearned models lack the knowledge necessary to empower malicious users (Figure 2).
For open-source models, unlearning can expunge hazardous knowledge before such models are publicly released, limiting adversaries from repurposing open-source models out of the box. However, unlearning on WMDP does not address the threat model of relearning in open source models We encourage future work towards mitigating risk in this pathway.
6.2 Structured API Access
@@
WMDP complements the safety benefits of structured API access (Shevlane, 2022), where model developers provide an API for users to query and finetune models without full weight access. In this framework, ordinary users may query and finetune models with an API, but the model provider applies safety mechanisms, such as unlearning, prior to serving the model. However, approved users could obtain API access to the base model with full capabilities under strict guidelines, empowering the use of LLMs for benign or defensive applications while mitigating potential vectors of malicious use. For instance, OpenAI allows access of GPT-4 variants with fewer guardrails for red-teaming and biological malicious use experiments (OpenAI, 2023a, 2024). Structured access mitigates the concern that unlearning dual-use information will harm defenders.
Structured access requires model developers to solve the “Know Your Customer” (KYC) challenge, which involves verifying the identity and intentions of customers before allowing them privileged interactions. For structured access, implementing KYC-like procedures can help mitigate the risks associated with malicious use by ensuring that only verified and trustworthy individuals or organizations are given the full capabilities of the model.
7 Conclusion
@@
We propose a dataset, WMDP, to evaluate the potential of malicious use in LLMs. WMDP was developed by subject matter experts in biology, cybersecurity, and chemistry, and was filtered to remove sensitive or export-controlled information. Modern LLMs score highly on some aspects of WMDP, suggesting presence of hazardous knowledge. We propose machine unlearning as a safety intervention to reduce hazardous knowledge.
Towards making progress on unlearning, we introduce RMU, an unlearning method that removes hazardous knowledge without significantly compromising general model performance. RMU also generalizes and successfully removes information from a private sensitive dataset. However, RMU reduces accuracy on closely related fields, such as introductory virology and computer security, demonstrating the need for continued research towards improved unlearning precision.
Acknowledgments